OpenStudy (anonymous):

Given: f(x) = x^3 - 4x What are all values of c that satisfy the conclusion of the mean value theorem on the interval [-3,3]

OpenStudy (anonymous):

First you want to find what the value of the function is at each of the boundaries: f(-3)=-15 f(3)=15 The slope of the secant line can be determined with this information, which is 5. Now we need to find all the values of c such that the derivative function is equal to 5: f'(c)=5=3c^2-4 9=3c^2c c=\[+/-\sqrt{3}\] We can check our answer by seeing what the slope is at c=+/-sqrt3: f'(sqrt3)=3(sqrt3)^2-4 =3(3)-4 =5 The same is true for -sqrt3. Therefore the two c values that prove the mean value theorem is x=+/-sqrt3

OpenStudy (anonymous):

so how did u get the slope 5?

OpenStudy (anonymous):

oh wait, nvm c:

OpenStudy (anonymous):

thank you <33

OpenStudy (anonymous):

can you help me with another one please? :)

OpenStudy (anonymous):

I determined slope by imagining a straight line that would go from (-3, -15) to (3, 15). If you take the change in y/change in x, you get 30/6, which is 5. But you got that already so good job =D

OpenStudy (anonymous):

OHH, i thought you just divided 15 by 3

OpenStudy (anonymous):

I can try, I was a little bit rusty with that mean value theorem problem, but I can try!